Stochastic Analysis

Place: Department of Mathematics, Kyoto University

Date : October 25 – 27, 2001

Program October 25 (Thur)

9:30 – 10:30 Itaru Mitoma (Saga Univ.) Asymptotic estimate of Chern-Simons integrals

10:40 – 11:40 Keisuke Hara (Ritsumeikan Univ.)

Quadratic Wiener functionals and dynamics on Grassmannians I

11:40 – 13:10 Lunch time

13:10 – 14:10 Tetsuya Kazumi (Osaka Prefecture Univ.)

Weyl transforms on infinite dimensional Gaussian spaces 14:20 – 15:20 Hideaki Uemura (Aichi Univ. of Education)

Unrenormalized intersection local time of Brownian motion and its local time representation

15:40 – 16:40 Isamu Dôku (Saitama Univ.)

Pathwise Regurality of Some Measure-Valued Processes 18:00 – 20:00 Dinner party at Kureshima October 26 (Fri)

9:30 – 10:30 Setsuo Taniguchi (Kyushu Univ.) Exponential decay of stochastic oscillatory integrals and its application

10:40 – 11:40 Keisuke Hara (Ritsumeikan Univ.)

Quadratic Wiener functionals and dynamics on Grassmannians II

11:40 – 13:10 Lunch time

13:10 – 14:10 Ichiro Shigekawa (Kyoto Univ.)

Boundary value problems in the Wiener space 14:20 – 15:20 Li-Ming Wu (Wuhan Univ.)

Spectral gap, uniform positive improvingness and tail norm. I 15:40 – 16:40 Li-Ming Wu (Wuhan Univ.)

Spectral gap, uniform positive improvingness and tail norm. II

October 27 (Sat)

9:30 – 10:30 Keisuke Hara (Ritsumeikan Univ.) Quadratic Wiener functionals and dynamics on Grassmannians III

10:40 – 11:40 Li-Ming Wu (Wuhan Univ.)

Spectral gap, uniform positive improvingness and tail norm. III

The talk of Kazuhiro Kuwae (Yokohama City Univ.) “Non-symmetric perturbations for symmetric Dirichlet forms” (Joint work with P.J.Fitzsimmons) is cancelled

Organizers: Ichiro Shigekawa (Kyoto Univ.)

Itaru Mitoma (Saga Univ.)